Suppose that for an isoquant, if L=3 and K=5, then Q= 10. Is it possible that...

  1. Suppose that for an isoquant, if L=3 and K=5, then Q= 10. Is it possible that L=3 and K=6 also produce Q= 10 in the economical relevant portion of the isoquant? Why or why not. What will

    MPL be equal to?  
Answer #1

Isoquant referes to Equal quantity ( Iso means equal and quant means quantity). it means that different combination of inputs namely capital and labor which produce same level of output. It is also called Equal product curve or production indifference curve.

Slope of Isoquant = MRTSLk

MRTSLK=MPL÷ MPk

shapeandcurvatureofIsoquantdetermineelasticity of  ​​​substitution. ConvexitydetermineMRTSLK.

Marginal Rate of Technical substitution (MRTS) is the rate at which one input is reduced to increase the use of another input in order to produce same level of output.

Answer:

So it is possible to produce same quantity of 10 by using different combination of inputs namely L=3 & K=5 and L=3 & K =6 in the economical relevant portion of the Isoquant. The marginal productivity of labor is equal to W/P where P is price and W is wage.

MPL=W/P

MPk=R/P where R is rental price on capital and p is price.